OpenStudy (anonymous):

find the variation constant in which y varies inversely as x, and the following condition exists. y=35 when x=1/7

OpenStudy (amistre64):

to say that y varies inversely as x tends to be code for division y = k/x ; solve for k xy = k; 35(1/7) = 5 .... k = 5 y = 5/x

OpenStudy (amistre64):

you could think of a long thin balloon with an x on one side and a y on the other; when you squeeze one end; the other end gets bigger ... that is an inverse relationship.

OpenStudy (amistre64):

the volume in the balloon would be represented by "k" :) perhaps

OpenStudy (anonymous):

thank u...